9. Vistaril is available in a vial labeled 40 mg in 1 mL. 0.1 g of Vistaril has been prescribed IM.

How many mL should you withdraw from the vial in preparing to administer the drug?

10. You have available an ampule of Nitrostat labeled 250 mg in 1 cc. 0.4 g of Nitrostat has

been prescribed by the physician IV. What volume of drug will you prepare?

11. The physician has prescribed 0.075 g of Clinoril po. The Clinoril is available in scored tablets

labeled 150 mg. How many scored tablets should you administer?

12. 20 mg of Vistaril has been prescribed IM. How many mL should be prepared using a vial

labeled 15 mg in 1 mL?

13. Lanoxicaps are available in 0.4 mg tablets. 0.3 mg has been prescribed po. You should

administer how many tablets?

14. Lasix is available in a vial labeled 300 mg per 2 mL. 0.2 g has been prescribed by the

physician po. You should prepare how many mL of the drug for administration?

15. 18 mg of liquid Dilantin has been prescribed po by the physician. What volume of Dilantin

should you administer if the Dilantin is available in a strength labeled 40 mg in 3 cc?

23. 0.6 g of Chloromycetin has been prescribed IV. What volume should be prepared using a vial

labeled 200 mg in 1 cc?

24. The physician has prescribed 0.4 g of Suprax. On hand you have Suprax labeled 200 mg. 0.4 g

of Suprax is equivalent to how many mg?

25. The physician has prescribed 2 mg of Dilaudid for administration IM. The Dilaudid is available

in a vial containing 4 mg in 1 cc. In preparing the drug for administration, what volume

should you withdraw from the vial?

26. The physician has prescribed 0.2 g of Nitrostat for administration IV. The Nitrostat is available

in an ampule containing 500 mg in 1 mL. In preparing the drug for administration, how many

mL should you withdraw from the ampule?

27. 200 mg scored tablets of Diabinese are available. How many scored tablets should be

administered po if 0.5 g of Diabinese is prescribed?

28. Ergotrate is available in an ampule labeled 0.1 mg in 1 cc. 0.2 mg of Erogtrate has been

prescribed IM. What volume should you withdraw from the ampule in preparing to

administer the drug?

29. You have a vial of Garamycin labeled 30 mg in 1 mL. 50 mg of Garamycin has been prescribed

by the physician IV. What volume of drug will you prepare?

30. 200 mg tablets of Lithium Carbonate are available. How many tablets should be

administered po if 0.5 g of Lithium Carbonate is prescribed?

31. You have on hand 20 mg of Vistaril. The physician has prescribed 0.2 g of Vistaril. 0.2 g of

Vistaril is equivalent to how many mg?

32. You have available a vial of Ancef labeled 220 mg in 1 mL. 300 mg of Ancef has been

prescribed by the physician IV. What volume of the drug will you prepare?
Answer #1

ANSWER.

9) Calculation of volume(ml) of drug required.

        Formula

           Volume(ml) = Desired dose/Dose in hand * Quantity.

   Here

        Desired dose(D) = 100 mg (0.1 g)

        Dose in hand (H) = 40 mg.

            Quantity (Q) = 1 ml

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 100/40 * 1

                                                   = 2.5 * 1

                                                   = 2.5 ml.

10) Calculation of volume(ml) of drug required.

            Volume(ml) = Desired dose(D) / Dose in hand (H) * Quantity (Q)

      Here

           Desired dose (D) = 400 mg (0.4 g)

           Dose in hand (H) = 250 mg.

               Quantity (Q) = 1 ml

   As per above formula

         Volume(ml) of drug required = D/H * Q

                                                       = 400/250 * 1

                                                       = 1.6 ml.

11) Calculation of number of tablet required.

           Ordered dose = 0.075 gm.

      Convert gm to mg for calculation

                1gm = 1000 mg.

       So

             0.075 gm = 0.075 * 1000

                              = 75 mg.

          Available dose = 150 mg

   Number of tablet required = 75 / 150

                                             = 0.5 ( Half of a tablet).

12) Calculation of volume(ml) of drug required.

            Volume(ml) = Desired dose(D)/Dose in hand (H) * Quantity(Q)

      Here

           Desired dose(D) = 20 mg.

           Dose in hand (H) = 15 mg.

               Quantity(Q) = 1 ml.

As per above formula

     Volume(ml) of drug required = D/H *Q

                                                    = 20/15 * 1

                                                    = 1.3 ml.

13) Calculation of number of tablets required.

              Ordered dose = 0.4 mg.

              Available dose = 0.3 mg/tablet.

       Number of tablets required = 0.4 / 0.3

                                                   = 1.33

                                                   = 1 tablet.

14) Calculation of volume(ml) of drug required.

          Volume(ml) = Desired dose(D)/Dose in hand (H) * Quantity(Q)

       Here

           Desired dose(D) = 200 mg (0.2 g)

           Dose in hand (H) = 300 mg.

               Quantity(Q) = 2 ml

   As per above formula

         Volume(ml) of drug required = D/H * Q

                                                        = 200/300 * 2

                                                        = 0.67 * 2

                                                        = 1.3 ml.

15) Calculation of volume of drug rewwquujddm.

           Volume(ml) = Desired dose(D)/Dose in hand(H) *Quantity(Q).

      Here

         Desired dose (D) = 18 mg.

         Dose in hand (H) = 40 mg.

             Quantity(Q) = 3 ml

As per above formula

     Volume(ml) of drug required = D/H *Q

                                                    = 18/40 * 3

                                                    = 0.45 * 3

                                                    = 1.4 ml.

23) Calculation of volume(ml) of drug required.

           Volume(ml) = Desired dose(D)/Dose in hand (H) * Quantity(Q)

      Here

          Desired dose(D) = 600 mg (0.6 g)

          Dose in hand (H) = 200 mg.

              Quantity(Q) = 1 ml.

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 600/200 * 1

                                                     = 3 * 1

                                                     = 3 ml.

24) Calculation of mg in 0.4 g of suprax.

                 1 g = 1000 mg

                0.4 g = X mg

        1 * X = 1000 * 0.4

              X = 400 mg.

     So

         0.4 g suprax in gm is 400 mg of suprax.

25) Calculation of volume(ml) of drug required.

            Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

      Here

          Desired dose(D) = 2 mg.

          Dose in hand (H) = 4 mg.

               Quantity(Q) = 1 ml.

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 2/4 * 1

                                                     = 0.5 * 1

                                                     = 0.5 ml.

26) Calculation of volume(ml) of drug required.

           Volume(ml) = Desired dose(D)/Dose in hand (H) * Quantity(Q)

     Here

           Desired dose(D) = 200 mg (0.2g)

           Dose in hand (H) = 500 mg.

               Quantity(Q) = 1 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                     = 200/500 * 1

                                                     = 0.4 ml.

27) Calculation of number of tablets required.

           Ordered dose = 500 mg (0.5 g)

           Available dose = 200 mg/tablet.

     Number of tablet required = Ordered dose/Available dose.

                                                = 500/200

                                                = 2 and half tablet.

28) Calculation of volume(ml) of drug required.

             Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

       Here

           Desired dose(D) = 0.2 mg.

           Dose in hand (H) = 0.1 mg.

                Quantity(Q) = 1 ml

    As per above formula

        Volume(ml) of drug required = D/H * Q

                                                       = 0.2/0.1 * 1

                                                       = 2 ml.

29) Calculation of volume(ml) of drug required.

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

     Here

         Desired dose(D) = 50 mg.

         Dose in hand (H) = 30 mg.

              Quantity(Q) = 1 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 50/30 * 1

                                                    = 1.7 ml.

30) Calculation of number of tablets required.

             Ordered dose = 500 mg (0.5g)

             Available dose = 200 mg.

      Number of tablets required = 500/200

                                                  = 2.5 ( 2 and half tablet).

31) 0.2 gm of vistaril in mg.

             1 gm = 1000 mg.

             0.2 gm = X mg.

        1 * X = 1000 * 0.2

             X = 200 mg.

   0.2 gm of vistaril in mg is 200 mg.

32) Calculation of volume(ml) of drug required.

             Volume(ml) = Desired dose(D) / Dose in hand(H) * Quantity(Q)

       Here

             Desired dose(D) = 300 mg.

             Dose in hand(H) = 220 mg.

                  Quantity(Q) = 1 ml.

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                    = 300/220 * 1

                                                    = 1.4 * 1

                                                   = 1.4 ml.